OpenStudy (anonymous):

When finding the sum of the arithmetic series 4 + 1 + –2 + –5 + –8 + –11, you will have a step of work that contains the number –3. Use complete sentences to describe how the –3 was calculated and what it means.

OpenStudy (anonymous):

We notice that the difference between a following term and its preceding term is -3, therefore, between each term, we recursively add -3 to reach the next term.

OpenStudy (campbell_st):

the 1st term in the sequence is a = 4 find the common difference, which means the difference between terms is always the same. \[a_{2} - a_{1} = a_{3} - a_{2} = a_{4} - a_{3} = ...\] there are 6 terms in your sequence.... there are several formulae that can be used \[S_{n} = \frac{n}{2}[ 2a + (n - 1)d]\] n = number of terms, a = 1st term, d = common difference A term in the series can be found using \[T_{n} = a + (n - 1)d\] so the above formula can be written as \[S_{n} = \frac{n}{2}[ a + a + (n - 1)d]\] or \[S_{n} = \frac{n}{2}[ a + l] \] l = last term a simple way to solve this problem is is reverse the order and add the 2 sets of numbers together 4, 1, -2, -5, -8, -11 + -11,-8, -5, -2, 1, 4 ------------------------- -7, -7, -7, -7, -7, -7 there are 6 lots of - 7 which sum to - 42 becuase you doubled the sets of numbers halve your answer -42/2 = -21 this is both formulae is a simple style.